# Step-by-Step Mastery: Enhancing Soft Constraint Following Ability of Large Language Models

### v2

@@ 3.3.1 Diversity
Complex to Simple He et al. (2024a) 1467 H $\checkmark$ $\checkmark$
Ours 7049 H/S $\checkmark$ $\checkmark$
Table 1: Detailed comparison of relevant works. Ours represents our dataset construction approach. 'Nums.', 'Cons.', 'Pair.', and 'Open.'  denote the number of samples, constraint types, whether to perform pairwise comparison, and whether the data is open-source.
@@ 3.3.2 Comparison with Other Works
As shown in Tab. 1, we compare our dataset construction approach with other related works. In comparison to existing methods, the scale of our dataset is large. From the perspective of constraint categories, our dataset includes both soft and hard constraints. Compared to datasets that only contain soft or hard constraints, this broader scope better facilitates the model’s ability to learn constraint following. Regarding pairwise comparison, we use Judger for comparisons of the outputs, which improves the quality of the dataset. Moreover, our dataset is open-source. To prevent catastrophic forgetting McCloskey and Cohen (1989) during training, we mix the data in each curriculum with a corresponding proportion of ShareGPT data Chiang et al. (2023).
Figure 3: Verb-noun structure of multi-constraint instructions.
@@ 4 Experiments
Qwen2-72B-Instruct Yang et al. (2024) Qwen 67.9 56.6 47.8 42.2 35.3 50.0 77.1 84.4 80.4 86.9 82.2
@@
Conifer-13B Sun et al. (2024) LLaMA2 60.5 53.6 48.4 40.7 31.7 47.0 42.9 53.0 47.5 57.4 50.2
Vicuna-13B-v1.5 Chiang et al. (2023) LLaMA2 71.2 60.2 49.6 40.6 34.0 51.1 43.1 53.6 46.6 58.0 50.3
Conifer-7B-SFT Sun et al. (2024) Mistral 54.3 49.5 49.3 40.8 30.5 44.9 45.8 57.1 50.8 62.0 53.9
Conifer-7B-DPO Sun et al. (2024) Mistral 60.3 53.6 48.0 47.1 41.0 50.0 48.1 59.1 52.3 63.3 55.7
@@
Mistral-7B-Instruct-v0.3SFT Mistral 58.7 52.4 42.5 37.2 35.6 45.3 56.8 67.8 60.6 71.3 64.1
Mistral-7B-Instruct-v0.3DPO+Judger+CL Mistral 61.8 56.8 44.9 39.8 33.8 47.4 53.4 63.5 57.1 67.5 60.4
@@
Llama-3-8B-InstructSFT LLaMA3 69.3 59.0 50.1 44.8 32.0 51.0 68.8 76.6 71.2 78.7 73.8
Llama-3-8B-InstructDPO+Judger+CL LLaMA3 69.2 59.6 50.8 48.9 44.6 54.6 72.5 80.3 77.1 84.1 78.5
Table 2: The overall performance on FollowBench and IFEval. We use boldface for the best results and underline for the second-best results among the models ranging from 7B to 13B parameter sizes. indicates that the results are directly sourced from the original benchmarks.
We conduct extensive experiments to evaluate the effectiveness of our proposed method, focusing on soft constraint following ability and generalization performance.
@@ 4.1 Experiment Setup
Models. We conduct experiments on two widely recognized base LLMs, Llama-3-8B-Instruct Dubey et al. (2024) and Mistral-7B-Instruct-v0.3 Jiang et al. (2023a), both of which demonstrate exceptional performance among models within the parameter range of 7B to 8B. Within our experimental framework (§3), we compare three approaches: (1) BASE directly utilizes the original model to generate outputs. (2) SFT applies supervised fine-tuning on LLMs using constructed data (§3.1.1). (3) DPO+Judger+CL utilizes Judger to produce high-quality training data, in accordance with training the model using DPO based on curriculum learning(§3.1.2, §3.2).
For baseline comparisons, we select a range of open-source and proprietary LLMs. Among the proprietary models, we include GPT-4 Achiam et al. (2023) and GPT-3.5-turbo. Additionally, we compare our approach with several open-source LLMs, including models specifically trained to improve general instruction-following abilities, such as Vicuna-13B-v1.5 Chiang et al. (2023). We also include models focused on enhancing the ability to follow soft instructions, such as WizardLM-v1.2-13B Xu et al. (2023) and the Conifer series Sun et al. (2024). We also compare our models against two 70B-sized models, Llama-3.1-70B-Instruct Dubey et al. (2024) and Qwen2-72B-Instruct Yang et al. (2024), which are among the most powerful models.
Evaluation Benchmarks. IFEval Zhou et al. (2023a) is a benchmark designed to assess the adherence to hard constraints. It defines 25 distinct types of verifiable instructions and generates approximately 500 prompts, each containing between 1 and 3 constraints. These hard constraints are explicit and unambiguous, enabling programmatic validation of compliance. FollowBench Jiang et al. (2023b), is a benchmark that evaluates the ability of models to follow both soft and hard constraints across multiple levels of granularity, offering a comprehensive assessment of instruction-following capabilities.
@@ 4.2 Main Results
As shown in Tab. 4, our method significantly enhances the model’s ability to follow soft constraints, even outperforming the capabilities of larger models. Furthermore, it also markedly improves the model’s hard constraint following ability, demonstrating its effectiveness across both types of constraints. Specifically, when the models are trained using the DPO+Judger+CL method, a significant performance improvement is observed across both benchmarks, particularly on IFEval. The model’s performance improvement is particularly significant on complex tasks, especially at the L4-L5 difficulty levels in FollowBench. Specifically, Mistral-7B-Instruct-v0.3 shows an average improvement of 4.3% at the L4-L5 difficulty levels.
In comparison to models designed to enhance the ability to follow complex instructions, our model demonstrates superior performance on both benchmarks. Specifically, although the performance of Mistral-7B-Instruct-v0.3 on Followench is lower than Conifer-7B-SFT, its performance surpasses the Conifer model on both benchmarks after training. Moreover, our training paradigm effectively enhances the instruction-following ability of LLMs, even when working with models of smaller parameter sizes. Specifically, compared with models in the 13B category, the performance of Llama-3-8B-Instruct is initially weaker than that of WizardLM-v1.2-13B on FollowBench. But after training, its performance surpasses the 13B model on both benchmarks.
After supervised fine-tuning on the constructed instruction-response pairs, the performance of the Llama-3-8B-Instruct model decreases on both benchmarks. This decline can be attributed to the fact that the Llama-3-8B-Instruct model incorporates various specialized training techniques during its initial training.
@@ 4.3 Generalization Experiments
Model BaseModel LC Win Rate
GPT-4-0613 GPT 30.2
GPT-3.5-Turbo-0613 GPT 22.4
@@
WizardLM-13B-v1.2 LLaMA2 14.5
Vicuna-13B-v1.5 LLaMA2 10.5
@@
Llama-3-8B-InstructBASE LLaMA3 21.6
Llama-3-8B-InstructDPO+Judger+CL LLaMA3 22.0
Table 3: Evaluation on the AlpacaEval2.0 for general LLM instruction-following ability. indicates that the results are directly sourced from the original leaderboards.
Besides the ability to follow soft constraints, we also assess the model’s general instruction following abilities on AlpacaEval Li et al. (2023b). To avoid the length bias that AlpacaEval may correlate with response lengths, we use the AlpacaEval 2.0 Zhao et al. (2024a) to evaluate the general instruction following.
In our evaluation process, we first perform supervised fine-tuning on the model, followed by DPO training using the proposed training paradigm. Specifically, we use precomputed outputs of GPT-4 Turbo on AlpacaEval as reference outputs and employ GPT-4o as evaluators. As shown in the Tab. 4.3, our method leads to a significant improvement in the model’s general instruction-following ability, outperforming both models of comparable parameter scales and even larger models.
@@ 4.4 Ablation Studies
In this section, we conduct ablation experiments to assess the impact of Judger, as described in §3.1.2, and the curriculum-based training paradigm, outlined in §3.2, on the model’s ability to follow instructions. The Llama-3-8B-Instruct model is used as the base model, and evaluations are conducted on the IFEval and FollowBench benchmarks.
@@
As shown in Tab. 4, using the constructed data directly for SFT without Judger adjustments underperforms the full method on both benchmarks, even resulting in a slight performance decline relative to the base model. It is evident that performance decreases significantly at the L4-L5 levels of FollowBench. This observation suggests that Judger plays a critical role in ranking responses to more challenging instructions. In contrast, the model trained with DPO outperforms the SFT baseline, especially on IFEval, further emphasizing the effectiveness of the DPO training approach over SFT in constraint following tasks.
Additionally, the results indicate that randomly organizing DPO training data leads to a decrease in performance. In contrast, our curriculum-based approach where training data is organized based on the number of constraints in the instructions leads to a significant improvement in the model’s ability to follow soft constraints, particularly those at higher difficulty levels in L4-L5 levels of FollowBench. While its performance on IFEval, which only includes hard constraints, slightly decreases. These findings strongly validate the necessity of Judger for constructing high-quality outpus and the proposed curriculum learning paradigm for enhancing the model’s ability to follow soft constraints.
@@ 4.5 Analysis
Figure 4: Results across various constraint categories in FollowBench.
@@ 4.5.1 Category Analysis
[…]
Ranking Method Kendall Tau Distance Position Consistency
@@
Table 5: Results on Judger’s effectiveness in aligning data with human preferences.
Figure 5: Results of the model on FollowBench and IFEval across different training stages in curriculum learning.
@@ 4.5.2 The Role of Judger
In this section, we investigate the factors contributing to the effectiveness of the Judger in constructing high-quality outputs. Judger ranks the outputs to better alignment with human preferences. To examine the underlying effectiveness of the Judger, we conduct an experiment designed to evaluate whether it facilitates this alignment.
Specifically, we randomly select 100 output sets from the construction process in §3.1.1, each containing 3 to 5 outputs. These outputs are manually annotated with the correct rankings, which serve as the reference standard for comparison. We evaluate the rankings in three distinct scenarios: (1) sequential rankings, (2) rankings adjusted by Judger, and (3) rankings annotated by human experts.
To assess the similarity between these rankings, we employ two complementary metrics. The first is the Kendall Tau distance Kendall (1938), a statistical measure that quantifies the number of discordant pairs between two sequences, thereby reflecting the extent of their relative order differences. In addition, we introduce the position consistency metric, which quantifies the proportion of elements that occupy the same relative positions across both rankings. This metric provides a direct evaluation of the alignment between rankings at each specific position. The results, presented in Tab. 5, demonstrate that the rankings adjusted by the Judger exhibit greater alignment with human-annotated rankings when compared to sequential rankings. This finding suggests that Judger enhances the quality of the training data by improving its consistency with human judgments, thus making the preference data more reliable for training.
@@ 4.5.3 The Role of Curriculum Learning
In this section, we analyze the effects of the curriculum-based training paradigm at different stages of the training process. Specifically, we examine the performance of Llama-3-8B-Instruct with the full method across three training stages, each corresponding to a different level of curriculum learning difficulty. Stage0 represents the base model, while Stage3 and Stage5 represent the stages where the model completes the easy curriculum and the hard curriculum, respectively.
As shown in Fig. 5, our proposed training paradigm progressively enhances the model’s instruction following capability across various training stages. Specifically, after easy curriculum learning, the model trained in Stage3 demonstrates superior performance compared to the base model across tasks L1-L3. In contrast, the model’s performance at L4-L5 in Stage3 is lower than Stage0. The possible reason is that Stage3 may not have adequately prepared for the complexity of L4-L5. The gap between these difficulty levels could have led to the initial performance drop. Subsequentially, when the model progresses to Stage5, after hard curriculum learning, performance improves significantly at the difficlut levels L4-L5. The results on IFEval further support this conclusion. Stage0 demonstrates the lowest average performance across all indicators. After curriculum learning, there is a significant improvement in the model’s performance on IFEval. By initially focusing on simpler preference learning and gradually progressing to more complex one, the model’s ability to follow constraints improves incrementally. This progression enables the model to achieve better performance on increasingly difficult instruction following tasks.

### v3

@@ 3.3.1 Data Statistics
Complex to Simple He et al. (2024a) 12939 H $\checkmark$ $\times$
Ours 17043 H/S $\checkmark$ $\checkmark$
Table 2: A detailed comparison of related works. Ours represents our dataset. “Nums.”, “Cons.”, “Pair.”, and “Prog.”  denote the number of preference pairs, constraint types, whether to perform pairwise comparison, and whether the dataset is progressively constructed.
@@ 3.3.2 Comparison with Other Works
As shown in Tab. 2, we compare our dataset with related works. Our dataset is large in scale compared to these methods. In terms of constraint categories, it includes both soft and hard constraints, enhancing the model’s ability to learn to follow different types of constraints. Additionally, we use Judger for pairwise comparisons of outputs, improving the overall quality of the dataset. Moreover, our dataset is progressively constructed.
@@ 4 Experiments
Qwen2-72B-Instruct Yang et al. (2024) Qwen 82.4 68.2 62.6 61.2 54.7 65.8 77.1 84.4 80.4 86.9 82.2
@@
Vicuna-13B-v1.5 Chiang et al. (2023) LLaMA2 56.2 42.9 32.3 32.1 24.6 37.6 43.1 53.6 46.6 58.0 50.3
Mistral-7B-ShareGPT Jiang et al. (2023a) Mistral 51.6 45.8 38.3 25.8 20.7 36.5 43.6 53.5 47.3 57.8 50.6
LLaMA3-8B-ShareGPT Dubey et al. (2024) LLaMA3 57.3 49.0 42.3 30.5 27.8 41.4 37.7 49.5 42.3 53.8 45.8
Conifer-7B Sun et al. (2024) Mistral 54.3 49.5 49.3 40.8 30.5 44.9 45.8 57.1 50.8 62.0 53.9
@@
Mistral-7B-Instruct-v0.3SFT+Judger Mistral 58.7 52.4 42.5 37.2 35.6 45.3 56.8 67.8 60.6 71.3 64.1
Mistral-7B-Instruct-v0.3DPO+Judger+CL Mistral 63.3 56.0 47.5 40.0 36.7 48.7 53.4 63.5 57.1 67.5 60.4
@@
LLaMA2-13B-Chat-HFSFT+Judger LLaMA2 53.6 43.0 41.2 34.9 25.8 39.7 32.3 44.7 43.1 55.0 43.8
LLaMA2-13B-Chat-HFDPO+Judger+CL LLaMA2 53.6 48.3 42.5 29.2 29.9 40.7 35.1 46.6 48.2 58.9 47.2
[…]
LLaMA3-8B-InstructSFT+Judger LLaMA3 66.3 55.4 50.1 49.7 39.8 52.3 70.4 77.8 73.2 80.1 75.4
LLaMA3-8B-InstructDPO+Judger+CL LLaMA3 69.2 59.6 50.8 48.9 44.6 54.6 72.5 80.3 77.1 84.1 78.5
Table 3: The overall performance on FollowBench and IFEval. We use bold for the best results and underlined for the second-best results among the models ranging from 7B to 13B parameter sizes. indicates that the results are directly sourced from the original benchmarks.
We conduct extensive experiments to validate the effectiveness of our method on improving LLMs’ soft constraint following ability.
@@ 4.1 Experiment Setup
Models. We conduct experiments on three widely recognized base LLMs, LLaMA3-8B-Instruct Dubey et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023a), and LLaMA2-13B-Chat-HF Touvron et al. (2023). Within our experimental framework, we compare three approaches: (1) BASE directly utilizes the base model to generate outputs. (2) SFT+Judger applies supervised fine-tuning on LLMs using the instruction-response pairs $(I_{n},O_{w_{n}})$ generated by Judger (§3.1.13.1.2). (3) DPO+Judger+CL utilizes Judger to obtain high-quality training data, which is then used for DPO training following the curriculum learning paradigm (§3.1.2, §3.2).
For baseline comparisons, we select various open-source and proprietary LLMs. Among the proprietary models, we include GPT-4 Achiam et al. (2023) and GPT-3.5 Turbo. Additionally, we compare our approach with several open-source LLMs, including models specifically trained to improve general instruction following abilities and models focusing on enhancing the ability to follow complex instructions. We also compare our models with two 70B-sized powerful models.
Settings.
For each seed instruction, we progressively add five constraints. In the curriculum setting, we combine curriculum 1 to 3 into a simple curriculum and curriculum 4 to 5 into a difficult curriculum. The model is first trained on the simple curriculum, and then on the difficult one.
Evaluation Benchmarks.
IFEval Zhou et al. (2023a) is a benchmark designed to evaluate the ability to follow hard constraints. It defines 25 distinct types of verifiable instructions, each containing between 1 and 3 constraints. FollowBench Jiang et al. (2023b) is a benchmark that evaluates the ability of models to follow both soft and hard constraints across multiple levels of granularity.
@@ 4.2 Main Results
As shown in Tab. 4, our method significantly enhances the model’s ability to follow soft constraints. Compared to the BASE method, there is a significant performance improvement across both benchmarks when three models are trained using the DPO+Judger+CL method, particularly on IFEval. The improvement is significant on difficult soft constraint following tasks, especially at the L5 difficulty level in FollowBench. Specifically, LLaMA-3-8B-Instruct shows an improvement of 5.5% at the L5 difficulty level.
After training with the DPO+Judger+CL method, within the range of models with 8B parameters, Mistral-7B-Instruct-v0.3 and LLaMA-3-8B-Instruct outperform other models on both benchmarks. Within the 13B model category, LLaMA2-13B-Chat-HF outperforms WizardLM-v1.2-13B and Vicuna-13B-v1.5 on FollowBench. Although its performance on IFEval, which only includes hard constraints, is lower than that of the other two models, our method demonstrates superior results on FollowBench, a more complex task that incorporates both soft and hard constraints.
After training with the SFT+Judger method, there is a decline in the performance of LLaMA2-13B-Chat-HF and Mistral-7B-Instruct-v0.3. This drop is attributed to the model’s integration of various specialized training techniques during its initial training phase. Although the SFT+Judger method performs better than DPO+Judger+CL on IFEval for the Mistral-7B-Instruct-v0.3 model, the method even performs worse than the BASE method on FollowBench, failing to effectively improve the model’s ability to follow soft constraints. Moreover, the performance of the SFT+Judger method is worse than the DPO+Judger+CL method on both benchmarks for the other two models.
@@ 4.3 Generalization Experiments
Model BaseModel AlpacaEval2.0 MT-Bench
GPT-4-0613 GPT 30.2 9.18
GPT-3.5-Turbo-0613 GPT 22.4 8.39
@@
WizardLM-13B-v1.2 LLaMA2 14.5 7.20
Vicuna-13B-v1.5 LLaMA2 10.5 6.57
@@
DPO+Judger+CL LLaMA3 22.0 6.80
Table 4: Results of the length control win rate of AlpacaEval2.0 and the score of MT-Bench. indicates that the results are directly sourced from the original leaderboards.
Besides the ability to follow soft constraints, we also assess the model’s general instruction following abilities on AlpacaEval2.0 Zhao et al. (2024a) and MT-Bench Zheng et al. (2023). We first perform SFT on the model, followed by DPO+Judger+CL. Specifically, we use precomputed outputs of GPT-4 Turbo on AlpacaEval as reference outputs and GPT-4o as evaluators. As shown in the Tab. 4.3, our method improves the model’s general instruction following ability on both banchmarks.
@@ 4.4 Ablation Studies
Curriculum Setting FollowBench (HSR) IFEval
L1 - L3 L4 - L5 Avg Avg
BASE 56.3 44.9 51.7 74.3
{C1}{C2}{C3}{C4}{C5} 57.4 37.9 49.6 76.7
{C1,C2,C3}{C4,C5} 59.9 46.8 54.6 78.5
Table 6: Results of different curriculum setting.
In this section, we conduct ablation experiments to study the impact of Judger and curriculum learning on LLaMA3-8B-Instruct. As shown in Tab. 5, directly using the constructed data for SFT without Judger reordering underperforms the SFT+Judger method on both benchmarks, even worse than the base model. The performance decreases significantly at the L4-L5 levels of FollowBench. This suggests that Judger plays a critical role in ranking the model’s responses to challenging instructions. The model trained with DPO outperforms the SFT baseline, especially on IFEval, emphasizing the effectiveness of DPO over SFT in constraint following tasks. Additionally, the curriculum learning training paradigm improves the model’s ability to follow constraints on both benchmarks, particularly those at higher difficulty levels (L4-L5) of FollowBench. This validates the necessity of curriculum learning paradigm for enhancing the model’s ability to follow soft constraints.
We also explore the impact of curriculum setting. As shown in Tab. 6, combining Curriculum 1 to Curriculum 3 into an easy curriculum and Curriculum 4 to Curriculum 5 into a difficult curriculum significantly improves model’s performance on both benchmarks compared to five separate curricula, especially on FollowBench. If the model is trained on five separate curricula, there is no clear boundary between task complexities. The model struggles to generalize from easier tasks to more difficult ones. Grouping similar tasks together with a clear progression in difficulty helps the model better focus on mastering each level of complexity, preventing overfitting to individual tasks.
@@ 4.5 Analysis
Figure 4: Results across various constraint categories.
@@ 4.5.1 Category Analysis
In this section, we analyze the model’s performance across different types of constraints. Specifically, we compare the performance of LLaMA3-8B-Instruct with the BASE and DPO+Judger+CL method on FollowBench. As shown in Fig. 4, our method significantly improves the model’s performance across different types of constraints. The most notable improvement is observed in the Mixed category , which is defined as a composition of multiple constraint categories to simulate complex real-world scenarios. Our method enhances the model’s performance by 10.6% in this category, suggesting a notable enhancement in the model’s ability to handle complex constraints in real-world scenarios. For categories that contain only soft constraints, our method improves the model’s performance by 6% in the Style category. However, in the category of situation constraints, the performance of the DPO+Judger+CL method is worse than that of BASE. The reason is that our constructed situation constraints focus on soft constraints, while in FollowBench, the situation constraints include both soft and hard constraints. Our method struggles with the hard constraints in the situation category, such as complex situation reasoning, leading to the decrease in performance.
4.5.2 The Role of Progressive Construction
In this section, we analyze the role of progressively constructing the dataset. We compare two methods: w/o progressive and w/ progressive. w/o progressive means replacing both the instruction field and the chosen field with the corresponding ones from samples that include all constraints in the progressively constructed DPO dataset. As shown in Tab. 7, our progressive construction method enhances the model’s performance on FollowBench, especially at the difficlut level L4. Introducing all constraints at once makes the training process unstable. By progressively adding constraints, the model can focus on learning to follow one constraint at a time and learn how to balance constraints effectively.
Method FollowBench (HSR)
L1 L2 L3 L4 L5 Avg
BASE 67.8 54.5 46.6 50.6 39.1 51.7
w/o progressive 69.8 56.7 50.8 42.4 47.6 53.5
w/ progressive 69.2 59.6 50.8 48.9 44.6 54.6
Table 7: Results of different construction methods.
Ranking Method Kendall Tau Coefficient Position Consistency
@@ 4.5.2 The Role of Progressive Construction
Table 8: Results on Judger’s effectiveness.
Figure 5: Results of the model across different training stages in curriculum learning.
@@ 4.5.3 The Role of Judger
In this section, we investigate the role of Judger in obtaining high-quality outputs. Specifically, we randomly select 100 seed instructions and rank the outputs at each step of progressive construction . We evaluate the rankings in three scenarios: (1) w/o Judger: directly use GPT-4o to generate outputs without reordering, (2) w/ Judger: use Judger to reorder the outputs, and (3) rankings annotated by human experts, which serve as the reference standard for comparison.
To assess the similarity between these rankings, we employ two metrics. The first is the Kendall Tau Coefficient Kendall (1938),which measures the correlation between two rankings by assessing the agreement in the order of paired items. Another metric is position consistency, which quantifies the proportion of elements that occupy the same relative positions. As shown in Tab. 8, the rankings adjusted by the Judger exhibit greater alignment with human-annotated rankings. This suggests that Judger enhances the consistency of outputs with human judgments, thereby improving their quality.
@@ 4.5.4 The Role of Curriculum Learning
We analyze the effects of the curriculum learning paradigm. Specifically, we compare the performance of LLaMA3-8B-Instruct with the DPO+Judger+CL method across three training stages. Stage0 represents the base model, while Stage3 and Stage5 represent stages where the model completes the easy curriculum and the hard curriculum, respectively.
As shown in Fig. 5, our training paradigm progressively enhances the model’s constraint following capability across various training stages. Specifically, after easy curriculum learning, the model trained in Stage3 shows superior performance compared to the base model across L1-L3. The model’s performance at L4-L5 in Stage3 is lower than Stage0. The reason is Stage3 has not adequately prepared for the complexity of L4-L5. The gap between these difficulty levels leads to the performance drop. Subsequentially, when the model progresses to Stage5, after hard curriculum learning, the average performance improves significantly at the difficlut levels L4-L5. The results on IFEval further support this conclusion. Stage0 has the lowest average performance across all indicators. After curriculum learning, there is a significant improvement in the model’s performance on IFEval. Although Stage 5 performs slightly worse than Stage 3 on IFEval, Stage 5 shows a significant performance improvement on FollowBench, indicating the enhancement in the model’s ability to follow soft constraints.